# Fast training and sampling of Restricted Boltzmann Machines

### v2

@@ 5.2.1 Parallel Tempering approaches
where again the index $t$ indicates that the Hamiltonian is evaluated using the parameters $\bm{\Theta}^{t}$. This move satisfies detailed balance with our target equilibrium distribution $p(\bm{x})=\exp(-\mathcal{H}(\bm{x}))/Z$, ensuring that the moves lead to the same equilibrium measure. The reason for using the trajectory is that PT algorithms work very well when a system undergoes a continuous phase transition. As discussed in the appendix G, adjusting the inverse temperature can lead to discontinuous transitions for which PT is not known to perform very well, while following the trajectory, the system undergoes several continuous transitions as discussed in Decelle et al. (2018); Bachtis et al. (2024).
Our PTT approach relies on the ease of sampling configurations from an initial RBM. In classical PCD training, the early “non-specialized” models thermalize quickly because they behave like Gaussian distributions. For RCM+PCD, independent configurations can be efficiently sampled from the low-rank RBM using a static Monte Carlo algorithm, as described in Section A.1 of the SI. The trajectory flow then significantly speeds up convergence to equilibrium, especially in the presence of a multimodal structure. The time interval between successive models is chosen to keep the acceptance probability for exchanges between neighboring models (Eq. 5) around 0.25 (we observed that the total time (number of AGS multiplied by number of machines) needed to thermalize a PTT chain did not depend on the acceptance fixed in the interval [0.2,0.5]. In practice, effective models can be saved during training by monitoring the acceptance rate between the last saved model and the current one. Pre-trained models require far fewer saved models, as many are naturally positioned near major phase transitions. The number of models used for each sampling process, along with a detailed description of the algorithm, is provided in the SI.
In Fig. 4–C1-3, we present the number of jumps between clusters as a function of the total number of AGS steps, where in PT-like schemes this refers to 1 (AGS step + one swap proposal) multiplied by the total number of models run in parallel. In all cases, PTT sampling outperforms the other methods, showing particularly remarkable results for HGD. Additionally, our method has the advantage of not requiring extra training like Stacked Tempering or simulating unnecessary temperatures.
Moreover, the equilibrium samples obtained from different models during sampling can be used to compute the log-likelihood, similar to the standard PT approach Krause et al. (2020). In Fig. 3–right we show that the quality of this estimation is comparable with the online trajectory AIS. Another desirable advantage of the PT algorithms is that one can easily check the thermalisation of the MCMC process by investigating the ergodicity of the random walk in models Banos et al. (2010). In addition, we evaluated other indicators of sample quality and potential overfitting. As detailed analysis is in Sect. H of the SI. Our method consistently provides high-quality samples across all datasets and, thanks to precise likelihood computation, can effectively detect overfitting.
@@ 6 Conclusions
In this work, we tackle the challenges of training high-quality equilibrium RBMs on highly structured datasets. We focused on improving the sampling process and accurately evaluating test log-likelihood throughout training. Our findings reveal that for datasets that are tightly clustered in their principal component analysis representation, introducing a pre-training phase can enhance training quality. This phase involves encoding the principal directions of the data into the model using a convex optimization approach. However, this pre-training method, that is specific to RBMs, appears to have limited benefits for less structured datasets or for long trainings, as chains get out of equilibrium, and should be combined in the future with optimized MCMC methods to achieve its maximum capacity.
Furthermore, we developed a new approach that exploits the progressive learning of features during RBM training to enhance evaluation and sampling. This approach enables online and precise computation of log-likelihood and leads to an efficient sampling strategy akin to the standard parallel tempering algorithm. Our results demonstrate that both our proposed methods outperform traditional algorithms without the need for simulating models at unnecessary temperatures or training additional models to facilitate sampling. Furthermore, both the trajectory AIS method for log-likelihood estimation and the trajectory parallel tempering algorithm are highly versatile, making them applicable to more complex energy-based models. Additionally, the low-rank RBM model can serve as an effective pre-initialization step for deeper architectures, and the trajectory PT algorithm is suitable for direct use in any EBM, regardless of complexity.
7 Code availability
The code and datasets are available at https://github.com/anonymousrbm/iclr-2025
@@ A.1 The low-rank RBM and its sampling procedure
$\bm{w}=\sum_{\alpha=1}^{d}w_{\alpha}\bar{\bm{u}}_{\alpha}\bm{u}_{\alpha}^{\top},\qquad\text{with}\qquad(\bm{u}_{\alpha},\bar{\bm{u}}_{\alpha})\in{\mathbb{R}}^{N_{v}}\times{\mathbb{R}}^{N_{h}},$ (6)
@@
Then, given a magnetization $\bm{m}$, we can compute the minimizer $\bm{\mu}^{*}(\bm{m})$ of $\phi(\mu)-\bm{m}^{T}\mu$ which is convex, using e.g. Newton method which converge really fast since we are in small dimension. Note that in practice we will obviously use finite estimates of $\phi$, assuming $N_{v}$ is large enough. As a result we get $\bm{\mu}^{*}(\bm{m})$ satisfying implicit equations given by the constraints given at given $N_{v}$:
@@
The training can therefore be done directly in the subspace of dimension $d$. In Ref. Decelle & Furtlehner (2021a), it has been shown that such RBM can be trained by mean of the Restricted Coulomb Machine, where the gradient is actually convex in the parameter’s space. It is then possible to do a mapping from the RCM to the RBM to recover the RBM’s parameters. In brief, the training of the low-dimensional RBM is performed by the RCM, and then the parameters are obtrained via a direct relation between the RCM and the RBM’s parameters. The detail of the definition and of the training of the RCM is detailed in the appendix A.2.
@@ A.2 The Restricted Coulomb Machine
As introduced in Decelle & Furtlehner (2021a), it is possible to exactly train a surrogate model for the RBM, called the Restricted Coulomb Machine (RCM), on a low dimensional dataset without explicitly sampling the machine allowing to learn even heavily clustered datasets.
@@
$m_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}v_{i}u_{i\alpha},\quad w_{\alpha a}\coloneqq\sum_{i=1}^{N_{v}}w_{ia}u_{i\alpha},\quad\theta_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}\theta_{i}u_{i\alpha}$ (20)
with $\alpha\in\{0,\ldots,d\}$ and $\bm{u}$ the projection matrix of the PCA.
@@
$p_{\text{RBM}}(\bm{m})=\frac{\exp\left(N_{v}\left[\mathcal{S}(\bm{m})+\sum_{\alpha=1}^{d}\theta_{\alpha}m_{\alpha}+\frac{1}{N_{v}}\sum_{a=1}^{N_{h}}\log\cosh\left(\sqrt{N_{v}}\sum_{\alpha=1}^{d}m_{\alpha}w_{\alpha a}+\eta_{a}\right)\right]\right)}{Z}$ (21)
@@
$p_{\text{RCM}}(\bm{m})=\frac{\exp\left(N_{v}\left[\mathcal{S}(\bm{m})+\sum_{\alpha=1}^{d}\theta_{\alpha}m_{\alpha}+\sum_{a=1}^{N_{h}}q_{a}\left|\sum_{\alpha=1}^{d}n_{\alpha}m_{\alpha}+z_{a}\right|\right]\right)}{Z}$ (23)
@@
$q_{a}=\sqrt{N_{v}\sum_{\alpha=1}^{d}w_{\alpha a}^{2}},\quad n_{a}=\frac{w_{\alpha a}}{\sqrt{\sum_{\alpha=1}^{d}w_{\alpha a}^{2}}},\quad z_{a}=\frac{\eta_{a}}{\sqrt{N_{v}\sum_{\alpha=1}^{d}w_{\alpha a}^{2}}}.$ (24)
@@
$w_{\alpha a}=\frac{1}{\sqrt{N_{v}}}q_{a}n_{a}\qquad\text{and}\qquad\eta_{a}=q_{a}z_{a},$
in order to obtain the RBM from the RCM. The model is then trained through log-likelihood maximization over its parameters. However, this objective is non-convex if all the parameters are trained through gradient ascent. To relax the problem, since we’re in low dimension, we can define a family of hyperplanes $(\bm{n},\bm{z})$ covering the space and let the model only learn the weights of each to the hyperplane. We can then discard the ones with a weight low enough for the approximation equation 22 to be bad.
@@ A.3 Learning the bias
To initialize the model with its bias set to the dataset’s mean, we define the 00-th direction of the longitudinal space as the direction obtained by projecting the configuration space onto the normalized mean vector of the dataset.
@@
$H(\bm{m}(v))=-\sum_{\alpha=0}^{N_{v}}\theta_{\alpha}m_{\alpha}-\frac{1}{N_{v}}\sum_{a}\log\cosh\left(\sqrt{N_{v}}\sum_{\alpha=1}^{d}m_{\alpha}w_{\alpha a}+\eta_{a}\right).$ (27)

### v3

@@ 4.2 Exploiting the training trajectory for sampling
where again the index $t$ indicates that the Hamiltonian is evaluated using the parameters $\bm{\Theta}^{t}$. This move satisfies detailed balance with our target equilibrium distribution $p(\bm{x})=\exp(-\mathcal{H}(\bm{x}))/Z$, ensuring that the moves lead to the same equilibrium measure. The reason for using the trajectory is that PT-like algorithms work very well when a system undergoes a continuous phase transition. As discussed in the SI G, adjusting the inverse temperature can lead to discontinuous transitions for which PT is not known to perform very well, while following the trajectory, the system undergoes several continuous transitions as discussed in (decelle2018thermodynamics; bachtis2024cascade). The pseudocodes for the standard PT and our new PTT algorithms are provided in SI B.1.
Figs. 3–A1-3 show the evolution of two independent Markov chains on the most trained model using the PTT sampling, showing a significant increase in the number of jumps compared to Figs. 3–B1-3. We also compare the number of mode jumps observed with PTT with those of other algorithms, as a function of the total number of AGS steps performed. For the PT algorithm, we test different values of $N_{T}$ and include the variant with a reference configuration (krause_algorithms_2020). For Stacked Tempering, we replicate the method of (roussel2023accelerated). The total AGS steps are taken as the product of the number of parallel models and the AGS steps per model, as shown in Fig. 3–C1-3. PTT consistently outperforms all other algorithms in all datasets.
Our PTT approach leverages the ease of sampling configurations from an initial RBM. During classical PCD training, early “non-specialized” models thermalize quickly, behaving like Gaussian distributions. The same is true for RBMs pre-trained from the low-rank RBM of decelle2021exact, as independent configurations for the first machine can be efficiently sampled by a static Monte Carlo algorithm (see SI A.1). The trajectory flow greatly accelerates equilibrium convergence, particularly in multimodal structures. The time interval between successive models is set to maintain an acceptance probability for exchanges (Eq. 3) around 0.25. Notably, the total time required to thermalize a PTT chain (AGS steps × number of machines) is unaffected by acceptance values within [0.2, 0.5] (see SI B.2 for details). In practice, effective models (i.e. the parameters of the model) can be saved during training by monitoring the acceptance rate between the last saved model and the current one. Pre-trained models require far fewer saved models, as many are naturally positioned near major phase transitions. The number of models used for each sampling process, along with a detailed description of the algorithm, is provided in the SI. PTT offers the advantage of requiring no extra training (unlike Stacked Tempering) or unnecessary temperature simulations and can be fully applied to other energy-based models. Instead, the equilibrium samples obtained from different models during sampling can be used to compute the log-likelihood, similar to the standard PT approach (krause_algorithms_2020). In Fig. 2–right we show that the quality of this estimation is comparable with the online trajectory AIS. Another desirable advantage of the PT algorithms is that one can easily check the thermalisation of the MCMC process by investigating the ergodicity of the random walk in models (banos2010nature), as shown in SI B.2. In addition, we evaluated other indicators of sample quality and potential overfitting in SI  I. Our method consistently provides high-quality samples across all datasets and, thanks to precise likelihood computation, can effectively detect overfitting.
On the failure of thermal algorithms– Thermal sampling algorithms (such as Simulated Annealing or PT) have difficulties in efficiently sampling models that describe structured data sets. This is mainly because in these types of models, remote modes can simply disappear or appear discontinuously when the temperature is slightly changed. This phenomenon is known in physics as a discontinuous or first-order phase transition. For a detailed introduction to this effect and an illustration of its occurrence when changing the temperature in an RBM toy model, see SI H. The presence of this or these temperature transitions can affect the performance of the PT algorithm, as a change between two neighboring temperatures is rarely accepted (and therefore much denser temperature ladders are required), but in multimodal models the problem can be even more harmful and difficult to detect: PT can even forget configurations in certain minority clusters while maintaining high swap acceptances. We illustrate this problem in Fig. 4-A, where we compare the equilibrium distribution $p(\bm{m})$ of a well-trained machine trained with the HGD dataset when sampled at two different temperatures $\beta=0.9$ and $\beta=1$. This marginal distribution can be efficiently calculated using the tethered MC algorithm bereux2023learning. We can see that from $\beta=1$ to $\beta=0.9$ the East Asian cluster (the upper right cluster) disappears from the measure. Since the jump to this cluster is strongly suppressed with AGS, as shown in Fig. 3-C2, simulating the system at higher temperatures does not help to find this cluster more easily, but prevents it from being found with a standard 1-temperature swap + 1AGS scheme (because it will be never nucleated at low temperatures). Indeed, in  4-B we show the distribution of samples generated at $\beta=1$ by a long PT sampling run (with all swap acceptances above 0.8) completely forgetting the Asian cluster. One might wonder if this is indeed the right balance between clusters learned by the model, so we checked it by performing a long ($10^{6}$ AGS steps) run initialized with these PT-generated samples, finding that the missing clusters gradually resettle. Instead, the PTT samples correctly identify all the clusters and the sampled distribution statistics remain unchanged when a normal AGS run is re-run from that initialisation.
5 The low-rank RBM pretraining
Figure 4: On the left: Panel A compares the marginal distribution $p(m)$ along the first two principal components for the pretrained+PCD RBM on the HGD dataset at inverse temperatures $\beta\!=\!1$ and $\beta\!=\!0.9$. Panel B presents sampling results at $\beta\!=\!1$ using PT, PTT, and $10^{6}$ AGS steps, initialized from configurations generated by both algorithms. Right panels: Panels C and D compare equilibrium samples from RBMs trained with PCD alone vs. PCD initialized on low-rank RBMs for MNIST01, Mickey, HGD, Ising2D, and CelebA. Scatter plots and histograms of projections onto the first two principal components show data (black) vs. generated samples (red). Panel E tracks log-likelihood evolution (train: solid, test: dashed) using online Tr-AIS. Results, averaged over 10 low-rank RBM initializations, show minimal variance, as indicated by narrow shaded regions.
In (decelle2021exact), it was shown that an RBM with only a few modes in its $\bm{W}$ can be trained precisely through a mapping to a Restricted Coulomb Machine and solving a convex optimization problem (see SI A.2). This approach enables training RBMs with simplified $\bm{W}$:
$\bm{W}=\sum_{\alpha=1}^{d}w_{\alpha}\bar{\bm{u}}_{\alpha}\bm{u}_{\alpha}^{\top},\qquad\text{with}\qquad(\bm{u}_{\alpha},\bar{\bm{u}}_{\alpha})\in{\mathbb{R}}^{{N_{\mathrm{v}}}}\times{\mathbb{R}}^{{N_{\mathrm{h}}}},$ (4)
where $\{w_{\alpha}\}$ are the singular values of the coupling matrix and the right singular vectors $\{\bm{u}_{\alpha}\}_{\alpha=1}^{d}$ correspond exactly to the first $d$ principal directions of the data set. Under this assumption, $p(\bm{v})$ is a function of only $d$ order parameters, given by the magnetizations along each of the $\boldsymbol{u}_{\alpha}$ components, i.e. $m_{\alpha}(\bm{v})=\bm{u}_{\alpha}\cdot\bm{v}/\sqrt{{N_{\mathrm{v}}}}$. In particular,
$H(\bm{v})=-\sum_{\alpha=0}^{d}\theta_{\alpha}m_{\alpha}-\sum_{a}\log\cosh\left(\sqrt{N_{\mathrm{v}}}\sum_{\alpha=1}^{d}\bar{u}_{\alpha a}w_{\alpha}m_{\alpha}+\eta_{a}\right)=\mathcal{H}(\bm{m}(\bm{v})),$ (5)
where $\bm{m}=\left(m_{1},\ldots,m_{\alpha}\right)$ and $\theta_{\alpha}$ is the projection of the visible bias onto the extended $\bm{u}$ matrix. To initialize the visible bias as proposed in (montavon2012deep), we add an additional direction $\bm{u}_{0}$ associated with the magnetization $m_{0}$ generated by the part of the visible bias that is not contained in the intrinsic space of $\bm{W}$. This is an improvement over (decelle2021exact) because it effectively adds an additional direction at minimal cost. We also found that the addition of the bias term is crucial to obtain reliable low-rank RBMs for image data. The optimal parameters are obtained by solving a convex optimization problem (SI A.2). This yields a probability distribution $p(\bm{m})$ in a lower-dimensional space than $p(\bm{v})$, which can be efficiently sampled using inverse transform sampling. However, discretizing the $\bm{m}$-space limits practicality to intrinsic dimensions $d\leq 4$ due to memory constraints. These low-rank RBMs replicate dataset statistics projected onto the first $d$ principal components and, despite their simplicity, can approximate the dataset, as demonstrated in Supplemental Figure 19 for the 5 datasets.
At the start of RBM training, the model encodes the strongest PCA components through multiple critical transitions (Decelle_2017; decelle2018thermodynamics; bachtis2024cascade). Pretraining with a low-rank construction bypasses these transitions, avoiding out-of-equilibrium effects from critical slowing down, which becomes prohibitive in highly clustered datasets. Once the main directions are incorporated, mixing times stabilize at much lower values, as barriers between clusters shrink when modes align properly bereux2023learning, making it a more reliable starting point for PCD training. However, since mixing times still increase over training, pretraining primarily mitigates poor initialization effects, while issues from insufficient MCMC steps may reappear later. Bypassing these transitions significantly improves model quality and data statistics reproduction, especially for short training times. In Fig. 4, we compare equilibrium samples from two RBMs trained on five datasets with identical settings (see SI C) but different strategies: one trained from scratch using standard PCD (tieleman2008training), the other using PCD after low-rank RBM pretraining. We also tested Jarzynski reweighting (carbone2024fast) (see SI E), but results were unstable on most datasets and are omitted. Across all datasets, pretraining improves sample quality and test log-likelihood. The pre-trained+PCD RBM balances modes effectively, as confirmed by PCA projections, and consistently achieves higher log-likelihood values using online Tr-AIS. Samples are drawn with the PTT algorithm, both described in Section 4.1. Scatter plots highlight the importance of pretraining for balanced models in clustered datasets, significantly improving test log-likelihood. However, for less clustered datasets like CelebA and full MNIST, pretraining offers little benefit, as splitting distributions into isolated modes is unnecessary (see Fig. 20 in SI J).
@@ 6 Conclusions
In this work, we address the challenges of training, evaluating, and sampling high-quality equilibrium Restricted Boltzmann Machines (RBMs) on highly structured datasets. Our primary focus is on improving the sampling process and accurately estimating the log-likelihood during training. We introduce a novel approach that leverages the progressive feature learning inherent in RBM training to enhance both evaluation and sampling. This method facilitates precise, online computation of log-likelihood and enables an efficient sampling strategy, overcoming the limitations of the standard parallel tempering algorithm. Additionally, our results indicate that for datasets with tightly clustered structures in their principal component analysis (PCA) representation, incorporating a pre-training phase can improve training outcomes. This pre-training involves encoding the principal directions of the data into the model through a convex optimization technique. However, this RBM-specific pre-training method shows limited advantages for less structured datasets or extended training durations, where chains deviate from equilibrium. In the latter case, the presented training approach should be combined in the future with optimized MCMC methods to achieve its maximum capacity.
Code availability
The code and datasets are available at https://github.com/DsysDML/fastrbm
Acknowledgments
Authors acknowledge financial support by the Comunidad de Madrid and the Complutense University of Madrid through the Atracción de Talento program (Refs. 2019-T1/TIC-13298 & Refs. 2023-5A/TIC-28934), the project PID2021-125506NA-I00 financed by the “Ministerio de Economía y Competitividad, Agencia Estatal de Investigación” (MICIU/AEI/10.13039/501100011033), the Fondo Europeo de Desarrollo Regional (FEDER, UE) and the French ANR grant Scalp (ANR-24-CE23-1320).
@@ A.1 The low-rank RBM and its sampling procedure
$\bm{W}=\sum_{\alpha=1}^{d}w_{\alpha}\bar{\bm{u}}_{\alpha}\bm{u}_{\alpha}^{\top},\qquad\text{with}\qquad(\bm{u}_{\alpha},\bar{\bm{u}}_{\alpha})\in{\mathbb{R}}^{{N_{\mathrm{v}}}}\times{\mathbb{R}}^{{N_{\mathrm{h}}}},$ (6)
@@
Then, given a magnetization $\bm{m}$, we can compute the minimizer $\bm{\mu}^{*}(\bm{m})$ of $\phi(\mu)-\bm{m}^{T}\mu$ which is convex, using e.g. Newton method which converge really fast since we are in small dimension. Note that in practice we will obviously use finite estimates of $\phi$, assuming $N_{\mathrm{v}}$} is large enough. As a result we get $\bm{\mu}^{*}(\bm{m})$ satisfying implicit equations given by the constraints given at given $N_{\mathrm{v}}$}:
@@
The training can therefore be done directly in the subspace of dimension $d$. In Ref. (decelle2021exact), it has been shown that such RBM can be trained by mean of the Restricted Coulomb Machine, where the gradient is actually convex in the parameter’s space. It is then possible to do a mapping from the RCM to the RBM to recover the RBM’s parameters. In brief, the training of the low-dimensional RBM is performed by the RCM, and then the parameters are obtrained via a direct relation between the RCM and the RBM’s parameters. The detail of the definition and of the training of the RCM is detailed in the appendix A.2.
@@ A.2 The Restricted Coulomb Machine
As introduced in (decelle2021exact), it is possible to precisely train a surrogate model for the RBM, called the Restricted Coulomb Machine (RCM), on a low dimensional dataset without explicitly sampling the machine allowing to learn even heavily clustered datasets.
@@
$m_{\alpha}\coloneqq\frac{1}{\sqrt{{N_{\mathrm{v}}}}}\sum_{i=1}^{{N_{\mathrm{v}}}}v_{i}u_{i\alpha},\quad w_{\alpha a}\coloneqq\sum_{i=1}^{{N_{\mathrm{v}}}}W_{ia}u_{i\alpha},\quad\theta_{\alpha}\coloneqq\frac{1}{\sqrt{{N_{\mathrm{v}}}}}\sum_{i=1}^{{N_{\mathrm{v}}}}\theta_{i}u_{i\alpha}$ (20)
with $\alpha\in\{1,\ldots,d\}$ and $\bm{u}$ the projection matrix of the PCA.
@@
$p_{\text{RBM}}(\bm{m})=\frac{\exp\left({N_{\mathrm{v}}}\left[\mathcal{S}(\bm{m})+\sum_{\alpha=1}^{d}\theta_{\alpha}m_{\alpha}+\frac{1}{{N_{\mathrm{v}}}}\sum_{a=1}^{{N_{\mathrm{h}}}}\log\cosh\left(\sqrt{{N_{\mathrm{v}}}}\sum_{\alpha=1}^{d}m_{\alpha}w_{\alpha a}+\eta_{a}\right)\right]\right)}{Z}$ (21)
@@
$p_{\text{RCM}}(\bm{m})=\frac{\exp\left({N_{\mathrm{v}}}\left[\mathcal{S}(\bm{m})+\sum_{\alpha=1}^{d}\theta_{\alpha}m_{\alpha}+\sum_{a=1}^{{N_{\mathrm{h}}}}q_{a}\left|\sum_{\alpha=1}^{d}n_{\alpha}m_{\alpha}+z_{a}\right|\right]\right)}{Z}$ (23)
@@
$q_{a}=\sqrt{{N_{\mathrm{v}}}\sum_{\alpha=1}^{d}w_{\alpha a}^{2}},\quad n_{a}=\frac{w_{\alpha a}}{\sqrt{\sum_{\alpha=1}^{d}w_{\alpha a}^{2}}},\quad z_{a}=\frac{\eta_{a}}{\sqrt{{N_{\mathrm{v}}}\sum_{\alpha=1}^{d}w_{\alpha a}^{2}}}.$ (24)
@@
$w_{\alpha a}=\frac{1}{\sqrt{{N_{\mathrm{v}}}}}q_{a}n_{a}\qquad\text{and}\qquad\eta_{a}=q_{a}z_{a},$
in order to obtain the RBM from the RCM. The model is then trained through log-likelihood maximization over its parameters. However, this objective is non-convex if all the parameters are trained through gradient ascent. To relax the problem, since we’re in low dimension, we can define a family of hyperplanes $(\bm{n},\bm{z})$ covering the space and let the model only learn the weights of each to the hyperplane. We can then discard the ones with a low weights to keep the approximation (equation 22) good enough.
@@ A.3 Learning the bias
To initialize the model with its bias set to the dataset’s mean, we define the 0-th direction of the longitudinal space as the direction obtained by projecting the configuration space onto the normalized mean vector of the dataset.
@@
$\mathcal{H}(\bm{m}(v))=-\sum_{\alpha=0}^{{N_{\mathrm{v}}}}\theta_{\alpha}m_{\alpha}-\frac{1}{{N_{\mathrm{v}}}}\sum_{a}\log\cosh\left(\sqrt{{N_{\mathrm{v}}}}\sum_{\alpha=1}^{d}m_{\alpha}w_{\alpha a}+\eta_{a}\right).$ (27)